OpenStudy (anonymous):

The amount of money that Andy will have at the end of each week, w, can be expressed at A(w) = 20w + 300. The amount of money that Sam will have at the end of each week, w, can be expressed at S(w) = 15w + 400. They have decided to combine their savings accounts. Write a function that expresses the total amount, T(w) they have in their savings account at the end of each week.

OpenStudy (anonymous):

A: T(W)=5W-100 B: T(W)=35W+700 C: T(W)=-5W+100 D: T(W)= 300W^2+12500W+120000

OpenStudy (andrewishere):

Ok so combine like terms. Only numbers paired with "w" can be added or subtracted from other numbers that are attached to "w." So 20w+15w=35w. Then 300+400=700. So the new equation of their combined savings accounts would be T(w)=35w+700, or choice answer B.

OpenStudy (andrewishere):

I hope this helped :)